OpenStudy (anonymous):

please help guys! solve, using the square root method: 5x^2=60. Round the solution to the nearest tenth

OpenStudy (jlastino):

Divide 60 by 5 Then find the square root of that... 60/5 = 12 I think the square root is between 3 and 4...if you are taking up calculus you can do linear apporximation

OpenStudy (anonymous):

it needs to be an x,y?

OpenStudy (jlastino):

no...you just need to find the square root of 12. try using a calculator (cheats xD)

OpenStudy (jlastino):

or make x1=9 y1=3 x2= 12 y2=? delta x = 3 Differentiate sqrt x dx/(2sqrt(x) 3/2sqrt9 3/2(3) 3/6 delta y=1/2 add that to 3 3+ .5 3.5

OpenStudy (anonymous):

so it would be 3.5,-3.5

OpenStudy (jlastino):

yup

OpenStudy (anonymous):

or just 3.5

OpenStudy (jlastino):

no it's both

OpenStudy (jlastino):

'cause if you square -3.5 you also get 12

OpenStudy (anonymous):

thankyou!

OpenStudy (jlastino):

np if you do use a a calcu you'll get 3.46410162 (if you round it to the nearest tenth's you get 3.5 also xD)